OpenStudy (anonymous):

Find a so that the graph of f(x)=loga(x) contains the point (1/7, -14). Give your answer to 2 decimal places. a=

OpenStudy (amistre64):

we are givenan x and a y value from the point, and are asked to find "a" which is just the base

OpenStudy (amistre64):

y = loga(x) , to solve for "a", change of base it maybe y = ln(x)/ln(a) ; multiply thru by ln(a), and divide out the y ln(a) = ln(x)/y ; take the "e" of each side e^(ln(a)) = e^(ln(x)/y) ; simplify a = e^(ln(x)/y)

myininaya (myininaya):

y=log_a(x)=lnx/lna -14=ln(1/7)/lna -14lna=ln(1/7) lna=-1/14*ln(1/7) e^ e^ a=e^{-1/14*ln(1/7)} a=e^{ln(1/7)^(-1/14)} a=(1/7)^(-1/14) is what i get but i'm just a beginner so i could be wrong

OpenStudy (anonymous):

yes! that was right (: thank you guys.

myininaya (myininaya):

omg i can do math thats so weird

OpenStudy (amistre64):

yours looks to be good :) with the numbers added in it appears to adjust more, but yep :)

myininaya (myininaya):

im sorry i think you guys like to call it maths

myininaya (myininaya):

is tht right? lol

OpenStudy (amistre64):

its like a cat; it doesnt matter what you call it, it wont come until its ready :)